# Thermodynamic and quantum fluctuations of horizon area

### v1

@@ II Quantum fluctuations of black hole entropy
Here we consider quantum fluctuations of the area of the black hole horizon using the gravitational canonically conjugate variables. In the case of the black hole, these are the horizon area $A$ and the gravitational coupling $K=1/(4G)$. These canonically conjugate variables were used in particular for the calculations of the transition rate from the black hole to the white hole with the same mass.Volovik2022 This rate is determined by the integral $\int_{C}A(K)dK$ over the tunneling trajectory $C$.
@@ IV Black hole fluctuations via ensemble of Planck scale ”atoms”
Such unusual variance of the number of Planckons $N$, which is not proportional to $N$, is the consequence of the Tsallis-Cirto non-extensive statisticsTsallisCirto2013 ; Tsallis2020 with $\delta=2$. This statistics describes the non-extensive composition law for entropy in the processes of splitting and merging of black holes: (see Refs.Volovik2025TC ; Manoharan2025 and references therein):
$\sqrt{S(N_{1}+N_{2})}=\sqrt{S(N_{1})}+\sqrt{S(N_{2})}\,$, (10)
This non-extensive composition law is in the origin of the black-hole thermodynamics.

### v2

@@ II Quantum fluctuations of black hole entropy
Here we consider quantum fluctuations of the area of the black hole horizon using the gravitational canonically conjugate variables. In the case of the black hole, these are the horizon area $A$ and the gravitational coupling $K=1/(4G)$. For the black hole with mass $M$ one has:
$\frac{\partial A}{\partial\tau}=\frac{\partial M}{\partial K}\,\,,\,\,\frac{\partial K}{\partial\tau}=-\frac{\partial M}{\partial A}\,$, (1)
where $\tau=1/T$ is the time in Euclidean metric.
These canonically conjugate variables were used in particular for the calculations of the transition rate from the black hole to the white hole with the same mass.Volovik2022 This rate is determined by the integral $\int_{C}A(K)dK$ over the tunneling trajectory $C$, which connects black and white holes.
@@ IV Black hole fluctuations via ensemble of Planck scale ”atoms”
Such unusual variance of the number of Planckons $N$, which is not proportional to $N$, is the consequence of the Tsallis-Cirto non-extensive statisticsTsallisCirto2013 ; Tsallis2020 with $\delta=2$. This statistics describes the non-extensive composition law for entropy in the processes of splitting and merging of black holes (see Refs.Volovik2025TC ; Manoharan2025 and references therein):
$\sqrt{S(N_{1}+N_{2})}=\sqrt{S(N_{1})}+\sqrt{S(N_{2})}\,.$ (11)
This non-extensive composition law is in the origin of the black-hole thermodynamics, which distinguishes the Tsallis-Cirto $\delta=2$ entropy from other possible types of generalized entropy.Odintsov2022a ; Odintsov2022b ; Odintsov2025